# Hierarchical Augmentation and Distillation for Class Incremental Audio-Visual Video Recognition

### 11 Jan 2024 v1

@@ 5.3.9 Compared with large-scale fully-supervised method
To illustrate the rationality of the joint-training (upper-bound) results in our model framework, we compare the fully supervised performance of our model with that of state-of-the-art large-scale transformer video model UniFormerV2 [78]. UniFormerV2 focuses only on the visual modality and neglecting the audio modality. However, we perform fully supervised training by leveraging both visual and audio modalities to establish the upper bound of fully supervised performance. To ensure a fair comparison with UniFormerV2, we unify the modality for our experiments. Specifically, we test UniFormerV2 on the AVK-100, AVK-200, and AVK-400 datasets. Due to limited computational resources, we focused on the model UniFormerV2-L/14@336 with Frame $32\times 3\times 2$, which achieved 89.7% on AVK-400, only 0.3% lower than the best-performing model UniFormerV2-L/14@336 with Frame $64\times 3\times 2$. On the other hand, we also implement our method on the same datasets, considering solely the visual modality (HAD-V) and both modalities (HAD). The results are presented in Table VIII. From Table VIII, it is evident that UniFormerV2, utilizing large-scale transformers as their backbone, significantly outperform our HAD-V method on the AVK-100, AVK-200, and AVK-400 datasets, with improvements of 9.1%, 10.6%, and 13.6%, respectively. Furthermore, our HAD considering both audio and visual modalities surpasses HAD-V focusing only on visual information. However, the performance of HAD on AVK-100, AVK-200, and AVK-400 still falls short of UniFormerV2 with massive model parameters, by 2.7%, 2.2%, and 4.9%, respectively. These findings affirm the validity of the upper bound of joint-training performance in Table II and demonstrate the benefit of integrating both audio and visual modalities. This also inspires us to focus on developing multi-modal large-scale transformer video models in the future.
TABLE VIII: Comparison of fully-supervised learning in AVK-100, AVK-200, and AVK-400.
Methods Visual Audio AVK-100 AVK-200 AVK-400
UniFormerV2 [78] 95.4 91.0 89.7
HAD-V (Ours) 86.3 80.4 76.1
HAD (Ours) 92.7 88.8 84.8

### 10 Apr 2024 v2

@@ 5.3.9 Compared with large-scale fully-supervised method
To illustrate the rationality of the joint-training (upper-bound) results in our model framework, we compare the fully supervised performance of our model with that of state-of-the-art large-scale transformer video model UniFormerV2 [79]. UniFormerV2 focuses only on the visual modality and neglecting the audio modality. However, we perform fully supervised training by leveraging both visual and audio modalities to establish the upper bound of fully supervised performance. To ensure a fair comparison with UniFormerV2, we unify the modality for our experiments. Specifically, we test UniFormerV2 on the AVK-400 dataset. Due to limited computational resources, we focus on the model UniFormerV2-L/14 with Frame $16\times 3\times 4$, we also implement our method on the same dataset, considering solely the visual modality (HAD-V) and both modalities (HAD). The results are presented in Table VIII. From Table VIII, it is evident that UniFormerV2, utilizing large-scale transformers as their backbone, significantly outperform our HAD-V method on the AVK-400 dataset, with improvements of 17.0%, respectively. Furthermore, our HAD considering both audio and visual modalities surpasses HAD-V focusing only on visual information. However, the performance of HAD on AVK-400 still falls short of UniFormerV2 with massive model parameters, by 8.3%, respectively. These findings affirm the validity of the upper bound of joint-training performance in Table II and demonstrate the benefit of integrating both audio and visual modalities. This also inspires us to focus on developing multi-modal large-scale transformer video models in the future.
TABLE VIII: Comparison of fully-supervised learning in AVK-400.
Methods Visual Audio AVK-400
UniFormerV2 [79] 93.1
HAD-V (Ours) 76.1
HAD (Ours) 84.8